# Calibrated Probabilistic Interpolation for GEDI Biomass

### v1

@@ Abstract
To resolve this, we introduce Attentive Neural Processes (ANPs), a probabilistic meta-learning framework that explicitly conditions predictions on local observation sets and geospatial foundation model embeddings. Unlike static ensembles, ANPs learn a flexible spatial covariance function, allowing uncertainty estimates to expand in complex landscapes and contract in homogeneous areas. We validate this approach across five distinct biomes ranging from Tropical Amazonian forests to Boreal and Alpine ecosystems, demonstrating that ANPs achieve competitive accuracy while maintaining near-ideal uncertainty calibration. We demonstrate the operational utility of the method through few-shot adaptation, where the model recovers most of the performance gap in cross-region transfer using minimal local data. This work provides a scalable, theoretically rigorous alternative to ensemble variance for continental scale earth observation.
@@
The method enables scalable, context-aware uncertainty quantification for large-scale mapping without the computational bottlenecks of geostatistics.
@@ 1 Introduction
Forest aboveground biomass (AGB) is important for carbon accounting, climate change mitigation, and conservation planning (Dubayah et al., 2022; Duncanson et al., 2022). Accurate and high resolution biomass maps enable monitoring of deforestation, forest degradation, and carbon sequestration which are essential for implementing conservation programs and climate targets. NASA’s Global Ecosystem Dynamics Investigation (GEDI) mission provides high quality lidar measurements of forest structure, including calibrated estimates of AGBD at 25m footprint resolution (Dubayah et al., 2022). However, GEDI samples sparsely, with footprints separated by approximately 600m along-track in criss-crossing orbital patterns. To support operational forest monitoring, these sparse observations are interpolated to create wall-to-wall maps.
@@
However, biomass point estimates alone are insufficient for decision making because uncertainty quantification is essential (Araza et al., 2022). Conservation planners need to know where estimates are reliable to prioritize field validation efforts and assess risk (Song et al., 2023). Carbon accounting requires quantifying uncertainty to meet reporting standards. Despite this need, most biomass mapping modeling work provide only limited uncertainty representations, typically reporting cross-validation RMSE or ensemble variance as proxies for predictive uncertainty (Li et al., 2020; Xu et al., 2024; Zurqani, 2025; Turton et al., 2022). An important but often overlooked distinction exists between prediction intervals, which is quantifying uncertainty about individual observations, and confidence intervals of ensemble predictions, which is measuring disagreement between models fit on different data samples. When biomass mapping studies report uncertainty, do these uncertainties actually reflect the data?
@@
This gap is noteworthy for GEDI-based biomass mapping due to the mission’s irregular sampling pattern, and more broadly for ecological modeling where sparse but clustered sampling is the norm. Standard machine learning methods learn global relationships between satellite features and biomass, treating all predictions equivalently regardless of their spatial relationship to training data. Intuitively, however, uncertainty should depend on context. Predictions far from GEDI footprints should be more uncertain than those nearby, and predictions in heterogeneous landscapes should be more uncertain than in homogeneous areas. Ensemble based methods do not naturally encode these spatial relationships. Recent work has applied probabilistic deep learning to GEDI data using deep ensembles to estimate raw waveforms (Lang et al., 2022). However, their approach focuses on the 1D signal processing task (deriving height from a waveform at a single location). Our work addresses the 2D spatial interpolation task (predicting biomass between tracks), where uncertainty is driven by spatial distance and landscape heterogeneity.
@@
In this work, we demonstrate that standard machine learning methods for biomass mapping face limitations: hyperparameter configurations that maximize predictive accuracy produce poorly calibrated uncertainty estimates, while well calibrated configurations sacrifice accuracy. We show that these limitations stem from the distinction between prediction intervals and ensemble variance, and that it persists across diverse forest biomes. We introduce an alternative approach based on Neural Processes (NPs) (Garnelo et al., 2018; Kim et al., 2019), a class of models that explicitly condition predictions on observed data points and naturally produce calibrated uncertainty estimates. Geostatistics via Gaussian Processes (Rasmussen and Williams, 2005) provides the theoretically pure framework for this problem, but it scales cubically with the number of data points, making it operationally difficult for GEDI-scale datasets. Neural Processes offer a way to capture the spirit of GP-style spatial conditioning in a scalable, amortized fashion. We combine NPs with embeddings from a remote sensing foundation model trained on Sentinel-1 and Sentinel-2 data (Feng et al., 2025), demonstrating that this architecture can effectively map GEDI L4A biomass estimates to wall-to-wall predictions with well-calibrated uncertainty. Neural Processes treat any set of context observations $(x_{c},y_{c})$ as conditioning information, learn a distribution over functions given those contexts, and make predictions at arbitrary target locations $x_{t}$, which is directly aligned with the GEDI mapping problem.
@@
We provide a systematic evaluation of uncertainty calibration for GEDI-based biomass mapping, revealing that standard ensemble methods (Random Forest, XGBoost) exhibit poor calibration despite being widely used in the literature. Then, we show that Attentive Neural Processes achieve parity or better accuracy while maintaining near-ideal calibration instead of the tradeoff observed in baseline methods. We validate these findings across five diverse regions spanning temperate (Maine, US), alpine (South Tyrol, Italy), boreal (Hokkaido, Japan), tropical lowlands (Guaviare, Colombia), and tropical montane (Tolima, Colombia) forests, demonstrating consistent performance across biomes. ANPs exhibit more graceful degradation under spatial extrapolation when trained on one region and tested on ecologically distinct areas. The model appropriately increases uncertainty, which is a useful property for operational deployment.
@@
Our findings have implications for operational biomass mapping. The poor calibration of standard methods suggests that reported uncertainties are potentially unreliable, potentially leading to misguided conservation decisions or inaccurate carbon accounting. Neural Processes provide a principled alternative that produces trustworthy uncertainty estimates at competitive computational efficiency.
@@ 2.1.2 Remote Sensing Foundation Model Embeddings
Rather than using hand crafted spectral indices or raw satellite imagery, we use embeddings from Tessera (Feng et al., 2025), a transformer-based remote sensing foundation model pretrained on Sentinel-1 (C-band SAR) and Sentinel-2 (multispectral optical) imagery at global scale. Tessera produces 128-dimensional embeddings at 10m spatial resolution that encode rich spectral-temporal-structural information about land surface characteristics.
@@ 2.2.2 Training Objective
The model is trained to maximize the evidence lower bound (ELBO):
@@ 2.4 Spatial Cross-Validation
Standard random train-test splits are inappropriate for spatial data due to spatial autocorrelation, which can inflate performance estimates if nearby observations appear in training and test sets (Roberts et al., 2017). We use buffered spatial cross-validation for rigorous evaluation.
@@ 3.1 Within-Region Performance
Conversely, configurations that achieve good calibration sacrifice accuracy. XGBoost with shallow trees (depth=1, n=50) attains Z-std of 1.038, which is near ideal, but Log $R^{2}$ drops to 0.689, representing a meaningful performance decrease relative to the best accuracy configuration. This suggests a tradeoff where ensemble-based methods cannot simultaneously maximize both objectives.
@@
3.2 Performance-Efficiency and Accuracy-Calibration Limitations
To investigate the relationship between predictive accuracy, uncertainty calibration, and computational cost, we performed hyperparameter sweeps for Random Forest (60 configurations spanning tree depths 1-20 and ensemble sizes 50-1000) and XGBoost (60 configurations with depths 1-20 and boosting rounds 50-1000). Each configuration was evaluated over 5 seeds on the Guaviare test set. Complete results are provided in  C.
[…]
Figure 1: Performance-efficiency and accuracy-calibration limitations in baseline methods. Each point represents a hyperparameter configuration evaluated across 5 random seeds. Size indicates training time, where larger is longer training. Dashed lines show Pareto frontiers. Dotted horizontal line indicates ideal calibration (Z-score std = 1.0).
@@ 3.2 Performance-Efficiency and Accuracy-Calibration Limitations
Practitioners using reasonable hyperparameters such as Random Forest with depth=8-10 and n=500 trees (balancing performance and speed) or XGBoost with depth=4-6 and n=100-200 rounds (balancing accuracy and calibration) obtain miscalibrated uncertainty estimates (Z-std 2-5). Random Forest can theoretically achieve better calibration with much deeper trees, but with exponential scaling in time. Even with unlimited computational resources, Random Forest’s bootstrap based uncertainty has known theoretical limitations for reliable inference (Wager and Athey, 2018; Mentch and Hooker, 2016), ensemble variance measures training variability rather than predictive uncertainty, and asymptotic guarantees require assumptions (e.g. subsampling rates, tree correlation structure) that may not hold in practice.
@@ 3.3 Multi-Region Validation
These multi region results establish three findings. The limitations observed in ensemble methods is not specific to tropical forests, and perhaps represents a limitation across diverse biomes. ANP’s calibration quality is robust. Z-std remains within 0.7-1.1 across all regions despite varying forest types, biomass ranges, topographic complexity, and seasonal dynamics. ANP exhibits an important safety property with appropriate uncertainty expansion in challenging regions, which ensemble methods lack.
@@ 3.4.1 Zero-Shot Spatial Extrapolation
Both models exhibit degraded accuracy when extrapolating, with the magnitude depending on ecological similarity between training and test regions (Figure 2, bottom row). Within-biome transfers show moderate performance loss. Training on Maine (temperate) and testing on South Tyrol (alpine/temperate transition) yields Log $R^{2}$ of 0.28 for ANP and 0.20 for XGBoost. Both models maintain positive predictive skill despite geographic and topographic differences. Cross-biome transfers between ecologically similar regions also work reasonably. South Tyrol-trained models achieve Log $R^{2}$ of 0.54 (ANP) and 0.53 (XGBoost) when tested on Hokkaido, likely reflecting structural similarities between alpine and boreal forests.
However, large ecological distances produce degradation or failure. Training on Maine (temperate) and testing on Tolima (tropical) results in Log $R^{2}$ -0.19 for ANP and -0.02 for XGBoost. The inverse transfer (Tolima to Maine) is even worse. Both models achieve strongly negative $R^{2}$ (ANP: -1.02, XGBoost: -3.05), indicating predictions worse than using the regional mean. These failures are unsurprising. Satellite feature-biomass relationships differ fundamentally between tropical evergreen forests and temperate deciduous forests, but the results reveal a distinction in how models respond to distributional shift.
@@
This pattern holds across the majority of off-diagonal pairs. ANP maintains higher coverage than XGBoost in all 12 cross-region transfers (Figure 2, top row). The model’s expanded uncertainty intervals, even if imperfectly calibrated under extreme distributional shift, provide a safer failure mode by signaling when predictions should not be trusted. This triggers the need for region specific data collection or model adaptation, rather than producing overconfident estimates that could mislead carbon accounting or conservation decisions.
These cross-region results establish that zero-shot transfer across major biome boundaries is generally ineffective for biomass mapping. Both ANP and XGBoost show substantial accuracy degradation when source and target regions differ ecologically. However, ANP’s ability to recognize distributional shift through expanded uncertainty (albeit imperfectly) provides a safer failure mode than XGBoost’s predictions. This raises a practical question. Can ANPs leverage limited target region data to enable adaptation where zero-shot transfer fails?
@@ 3.5 Calibration Analysis
The quantile-quantile plot (Figure 4, bottom right) compares empirical quantiles of standardized residuals to theoretical quantiles of N(0,1). Points lying on the diagonal indicate perfect agreement. ANP shows excellent calibration in the central quantiles (approximately 2<z<2-2<z<2, covering  95% of predictions), with points closely following the diagonal line. Deviation appears primarily in the extreme tails where the empirical distribution has slightly heavier left tail (more extreme negative residuals) and thinner right tail (fewer extreme positive residuals) than the theoretical Gaussian. This asymmetry may reflect limitations in predicting the highest biomass values, consistent with known sensor saturation effects (Joshi et al., 2017). However, the tail deviations are minor and affect only the outermost 2-3% of predictions, while the bulk of the distribution (97% of cases) shows strong agreement with the Gaussian assumption.
@@ 3.6 Spatial Structure of Uncertainty
4 Discussions
@@ 4.3 Implications for Operational Mapping
Additionally, our results reveal a concerning pattern. The default hyperparameters commonly used in ecological and remote sensing applications produce systematically miscalibrated uncertainty estimates. Random Forest with 100-500 trees and moderate depth (5-10) achieves Z-score standard deviation of 4.5-7.0 with coverage of 30-45% (Table 7), far below the nominal 68% expected. Similarly, XGBoost with typical hyperparameters (100-200 rounds, depth 4-6) used for quantile regression produces either good accuracy with poor calibration or good calibration with degraded accuracy.

### v2

@@ Abstract
To resolve this, we introduce Attentive Neural Processes (ANPs), a probabilistic meta-learning architecture that explicitly conditions predictions on local observation sets and geospatial foundation model embeddings. Unlike static ensembles, ANPs learn a flexible spatial covariance function, allowing uncertainty estimates to expand in complex landscapes and contract in homogeneous areas. We validate this approach across five distinct biomes ranging from tropical Amazonian forests to boreal, temperate, and alpine ecosystems, demonstrating that ANPs achieve competitive accuracy while maintaining near-ideal uncertainty calibration. We demonstrate the operational utility of the method through few-shot adaptation, where the model recovers most of the performance gap in cross-region transfer using minimal local data. This work provides a scalable, theoretically rigorous alternative to ensemble variance for continental scale earth observation.
@@
The method enables scalable, context-aware uncertainty quantification for large-scale mapping.
@@ 1 Introduction
Forest aboveground biomass density (AGBD) is important for carbon accounting, climate change mitigation, and conservation planning (Dubayah et al., 2022; Duncanson et al., 2022). Accurate and high resolution biomass maps enable monitoring of deforestation, forest degradation, and carbon sequestration which are essential for implementing conservation programs and climate targets. NASA’s Global Ecosystem Dynamics Investigation (GEDI) mission provides high quality lidar measurements of forest structure, including calibrated estimates of AGBD at 25m footprint resolution (Dubayah et al., 2022). However, GEDI samples sparsely, with footprints separated by approximately 600m along-track in criss-crossing orbital patterns. To support operational forest monitoring, these sparse observations are interpolated to create wall-to-wall maps.
@@
However, biomass point estimates alone are insufficient for decision making because uncertainty quantification is essential (Araza et al., 2022). Conservation planners need to know where estimates are reliable to prioritize field validation efforts and assess risk (Song et al., 2023). Carbon accounting requires quantifying uncertainty to meet reporting standards. Despite this need, most biomass mapping studies provide only limited uncertainty representations, typically reporting cross-validation RMSE or ensemble variance as proxies for predictive uncertainty (Li et al., 2020; Xu et al., 2024; Zurqani, 2025; Turton et al., 2022). To our knowledge, no prior work has systematically evaluated whether reported uncertainties for GEDI-based machine learning biomass mapping are actually calibrated, or proposed methods specifically designed to produce reliable prediction intervals. An important but often overlooked distinction exists between prediction intervals, which quantify uncertainty about individual observations, and ensemble variance, which measures disagreement between models fit on different data samples. When biomass mapping studies report uncertainty, do these uncertainties actually reflect the data?
@@
This gap is noteworthy for GEDI-based biomass mapping due to the mission’s irregular sampling pattern, and more broadly for ecological modeling where sparse but clustered sampling is the norm. Recent work has validated prediction interval coverage for GLM-based biomass estimation using airborne LiDAR (Mukhopadhyay et al., 2024), but equivalent scrutiny has not been applied to the ML ensemble methods standard in GEDI wall-to-wall mapping. Standard machine learning methods learn global relationships between satellite features and biomass, treating all predictions equivalently regardless of their spatial relationship to training data. Intuitively, however, uncertainty should depend on context. Predictions far from GEDI footprints should be more uncertain than those nearby, and predictions in heterogeneous landscapes should be more uncertain than in homogeneous areas. Ensemble based methods do not naturally encode these spatial relationships. Recent work has applied probabilistic deep learning to GEDI data using deep ensembles to estimate raw waveforms (Lang et al., 2022). However, their approach focuses on the 1D signal processing task (deriving height from a waveform at a single location). Our work addresses the 2D spatial interpolation task (predicting biomass between tracks), where uncertainty is driven by spatial distance and landscape heterogeneity.
In this work, we demonstrate that standard machine learning methods for biomass mapping face fundamental limitations in uncertainty quantification, and that these limitations persist across diverse forest biomes. The core issue is a conflation of two distinct statistical quantities. Ensemble variance, what Random Forest typically report as uncertainty, measures model disagreement across bootstrap samples or random initializations. Prediction intervals measure something different, namely the range within which future observations are likely to fall. These answer different questions, and the former systematically underestimates the latter when models are well-tuned for accuracy (Section 4.1). Compounding this, standard methods treat predictions as independent given input features, with no mechanism to adapt uncertainty to local spatial context. A prediction at a forest-field boundary and a prediction in contiguous forest receive identical uncertainty estimates if their spectral signatures match, despite vastly different local variability (Section 4.2).
@@
We introduce an alternative approach based on Neural Processes (NPs) (Garnelo et al., 2018; Kim et al., 2019), a class of models that explicitly condition predictions on observed data points and naturally produce calibrated uncertainty estimates. Geostatistics via Gaussian Processes (Rasmussen and Williams, 2005) provides the theoretically pure framework for this problem, but it scales cubically with the number of data points, making it operationally difficult for GEDI-scale datasets. Neural Processes offer a way to capture the spirit of GP-style spatial conditioning in a scalable, amortized fashion. We combine NPs with embeddings from a remote sensing foundation model trained on Sentinel-1 and Sentinel-2 data (Feng et al., 2025), demonstrating that this architecture can effectively map GEDI L4A biomass estimates to wall-to-wall predictions with well-calibrated uncertainty. Neural Processes treat any set of context observations $(x_{c},y_{c})$ as conditioning information, learn a distribution over functions given those contexts, and make predictions at arbitrary target locations $x_{t}$, which is directly aligned with the GEDI mapping problem.
@@
We provide a systematic evaluation of uncertainty calibration for GEDI-based biomass mapping, revealing that standard ensemble methods (Random Forest, XGBoost) exhibit poor calibration despite being widely used in the literature. Then, we show that Attentive Neural Processes achieve parity or better accuracy while maintaining near-ideal calibration. We validate these findings across five diverse regions spanning temperate (Maine, US), alpine (South Tyrol, Italy), boreal (Hokkaido, Japan), tropical lowlands (Guaviare, Colombia), and tropical montane (Tolima, Colombia) forests, demonstrating consistent performance across biomes. ANPs exhibit more graceful degradation under spatial extrapolation when trained on one region and tested on ecologically distinct areas. The model appropriately increases uncertainty, which is a useful property for operational deployment. Finally, we demonstrate that ANPs enable few-shot adaptation to new biomes, recovering functional accuracy with minimal local data, a capability that tree-based ensembles lack due to their non-differentiable structure.
@@
Our findings have implications for operational biomass mapping. The poor calibration of standard methods suggests that reported uncertainties may be unreliable, potentially leading to misguided conservation decisions or inaccurate carbon accounting. Neural Processes provide a principled alternative that produces trustworthy uncertainty estimates.
@@ 2.1.2 Remote Sensing Foundation Model Embeddings
Rather than using hand crafted spectral indices or raw satellite imagery, we use embeddings from Tessera (Feng et al., 2025), a transformer-based remote sensing foundation model pretrained on Sentinel-1 (C-band SAR) and Sentinel-2 (multispectral optical) imagery at global scale. We retrieve these embeddings via the geotessera Python package and use them as frozen features; Tessera itself is not fine-tuned. Tessera produces 128-dimensional embeddings at 10m spatial resolution that encode rich spectral-temporal-structural information about land surface characteristics.
@@ 2.2.2 Training Objective
Neural Processes are trained using variational inference, which learns an approximate distribution over functions by optimizing a tractable lower bound on the data likelihood. The model is trained to maximize the evidence lower bound (ELBO):
@@ 2.4 Spatial Cross-Validation
Standard random train-test splits are inappropriate for spatial data due to spatial autocorrelation, which can inflate performance estimates if nearby observations appear in training and test sets (Roberts et al., 2017; Ploton et al., 2020). We use buffered spatial cross-validation for rigorous evaluation.
@@ 3.1 Within-Region Performance
3.2 Comparison to Hyperparameter-tuned Baselines
To investigate the relationship between predictive accuracy, uncertainty calibration, and model configurations, we performed hyperparameter sweeps for Random Forest (60 configurations spanning tree depths 1-20 and ensemble sizes 50-1000) and XGBoost (60 configurations with depths 1-20 and boosting rounds 50-1000). Each configuration was evaluated over 5 seeds on the Guaviare test set. Complete results are provided in C.
Figure 1: Performance in baseline methods. Each point represents a hyperparameter configuration evaluated across 5 random seeds. Horizontal line indicates ideal calibration (Z-score std = 1.0). Left: Random Forest improves monotonically but asymptotes above ideal calibration. Right: XGBoost shows scattered performance.
@@ 3.2 Comparison to Hyperparameter-tuned Baselines
[…]
@@ 3.4.1 Zero-Shot Spatial Extrapolation
Both models exhibit degraded accuracy when extrapolating, with the magnitude depending on ecological similarity between training and test regions (Figure 2, bottom row). Within-biome transfers show moderate performance loss. Training on Maine (temperate) and testing on South Tyrol (alpine/temperate transition) yields Log $R^{2}$ of 0.28 for ANP and 0.20 for XGBoost. Both models maintain positive predictive accuracy despite geographic and topographic differences. Cross-biome transfers between ecologically similar regions also work reasonably. South Tyrol-trained models achieve Log $R^{2}$ of 0.54 (ANP) and 0.53 (XGBoost) when tested on Hokkaido, likely reflecting structural similarities between alpine and boreal forests.
However, large ecological distances produce degradation or failure. Training on Maine (temperate) and testing on Tolima (tropical) results in Log $R^{2}$ -0.19 for ANP and -0.02 for XGBoost. The inverse transfer (Tolima to Maine) is worse. Both models achieve strongly negative $R^{2}$ (ANP: -1.02, XGBoost: -3.05), indicating predictions worse than using the regional mean. These failures are unsurprising. Satellite feature-biomass relationships differ fundamentally between tropical evergreen forests and temperate deciduous forests.
@@
These cross-region results establish that zero-shot transfer across major biome boundaries is generally ineffective for biomass mapping. Both ANP and XGBoost show substantial accuracy degradation when source and target regions differ ecologically. However, ANP’s ability to recognize distributional shift through expanded uncertainty (albeit imperfectly) provides a better performing failure mode than XGBoost’s predictions. This raises a practical question. Can ANPs leverage limited target region data to enable adaptation where zero-shot transfer fails?
@@ 3.5 Calibration Analysis
The quantile-quantile plot (Figure 4, bottom right) compares empirical quantiles of standardized residuals to theoretical quantiles of N(0,1). Points lying on the diagonal indicate perfect agreement. ANP shows excellent calibration in the central quantiles (approximately 2<z<2-2<z<2, covering  95% of predictions), with points closely following the diagonal line. Deviation appears primarily in the extreme tails where the empirical distribution has slightly heavier left tail (more extreme negative residuals) and thinner right tail (fewer extreme positive residuals) than the theoretical Gaussian. The tail deviations are minor and affect only the outermost 2-3% of predictions, while the bulk of the distribution (97% of cases) shows strong agreement with the Gaussian assumption.
@@ 3.6 Spatial Structure of Uncertainty
To quantify the relationship between observation density and predicted uncertainty, we computed block-level correlations for a representative tile. Dividing the tile into $\sim$1 km blocks (0.01), we calculated GEDI shot density and mean predicted uncertainty per block. Predicted uncertainty correlates significantly with observation sparsity (Pearson $r=0.43$, $p<10^{-5}$). Blocks with fewer GEDI shots show higher predicted uncertainty, confirming that the attention mechanism appropriately expands intervals in data-sparse areas.
4 Discussion
@@ 4.3 Implications for Operational Mapping
Additionally, our results reveal a pattern. The default hyperparameters commonly used in ecological and remote sensing applications produce systematically miscalibrated uncertainty estimates. Random Forest with 100-500 trees and moderate depth (5-10) achieves Z-score standard deviation of 4.5-7.0 with coverage of 30-45% (Table 7), far below the nominal 68% expected. Similarly, XGBoost with typical hyperparameters (100-200 rounds, depth 4-6) used for quantile regression produces either good accuracy with poor calibration or good calibration with degraded accuracy.
@@ 5.3 Active Learning
5.4 Propagating Observation-Level Uncertainty
The current framework treats GEDI L4A estimates as point observations. By treating all context observations as equally reliable, the model conflates two distinct sources of uncertainty in its aleatoric term in measurement noise (reducible with better instruments or repeated observations) and spatial heterogeneity (irreducible given the feature set). A natural extension would propagate observation-level uncertainties through the architecture. Several approaches are possible. Uncertainty-aware attention could downweight unreliable context points during aggregation, so that GEDI shots with high reported uncertainty contribute less to the predictive distribution. Alternatively, distributional context encoding could represent each observation as a distribution rather than a point, with the encoder learning that high-variance observations provide weaker constraints. Most directly, the latent prior could be conditioned on aggregate context uncertainty, widening $p(z|C)$ when the available observations are themselves uncertain.
This extension would enable the model to distinguish between predictions that are uncertain because nearby GEDI shots are unreliable versus predictions that are uncertain because the landscape is heterogeneous. Such disentanglement could improve interpretability for operational users and provide more principled uncertainty estimates when GEDI sampling quality varies across a study area. However, implementation requires careful consideration of how GEDI uncertainties are themselves derived and calibrated, as propagating miscalibrated input uncertainties could compound rather than correct biases in the final predictions.